OpenStudy (anonymous):

f(x) = {x^2-x / (x+1) if x does not = -1 { 0 if x = -1 Is this graph continuous at x = -1? Please explain! Thanks

OpenStudy (lgbasallote):

if x = -1 then what is f(-1)?

OpenStudy (helder_edwin):

remember the definition: f is continous at x=-1 if \[ \large \lim_{x\to-1}f(x)=f(-1) \]

OpenStudy (anonymous):

this is not as hard as it looks replace \(x\) by -1 in the expression \(\frac{x^2-x}{x+1}\) you will get \(\frac{2}{0}\) which is undefined so forget about it being continuous at -1\ if you got \(\frac{0}{0}\) you would have more work to do, but in this case there is nothing to compute after that